# Credit Risk Meets Large Language Models: Building a Risk Indicator from Loan Descriptions in P2P Lending

### v1

@@ 4. Description of the dataset
revenue
Borrower’s self-declared annual income during registration.
dti_n
Indebtedness ratio for obligations excluding mortgage. Monthly information.
loan_amnt
Amount of credit requested by the borrower.
fico_n
Credit bureau score. Defined between 300 and 850, reported by Fair Isaac Corporation as a summary risk measure based on historical credit information reported at the time of application.
@@
emp_length
Employment length of the borrower categorized into 12 categories, including the no information category.
purpose
Credit purpose category for the loan request.
home_ownership
Homeownership status provided by the borrower.
addr_state
Borrower’s residence state from the USA.
@@
desc
Description of the credit request provided by the borrower.
Regarding the variables that represent the features of the loans, we will consider that we are developing a granting model and, consequently, we will only use those variables known at the time of application. Table 1 shows the descriptions of the variables and we explain them below.
As for the quantitative variables, the information from the Fair Isaac Corporation credit bureau (FICO) in the original dataset is given by a minimum and maximum range of limits to which the borrower’s FICO belongs at loan origination. However, we average these two values to have a single indicator of the creditworthiness of potential borrowers resulting in our fico_n variable. For the case of the debt variable, dti_n is estimated from the original dataset variables as the ratio calculated from the total debts of the co-borrowers over the total debt obligation divided by the combined monthly income of the co-borrowers.
Regarding the categorical variables, we opted to merge the categories ‘other’, ‘none’, and ‘any’ into a unified category labeled ‘other’ for the home_ownership variable. This decision was made due to a lack of clear differentiation among these options, coupled with their similar default percentages and their relatively low percentages of occurrences. The emp_length variable was treated as categorical rather than numerical since it includes categories for ‘no information’ and for ‘more than ten years’.
For the textual variable, we carried out an exhaustive work of text cleaning. First of all, we removed all those descriptions that contained the default description provided by Lending Club on its web form (“Tell your story. What is your loan for?”). Moreover, we removed the prefix “Borrower added on DD/MM/YYYY $>$ from the descriptions, as we did not want any temporal background on them. Finally, as these descriptions came from a web form, we substituted all the HTML elements by their character (e.g. ‘&amp;’ was substituted by ‘&’, ‘&lt;’ was substituted by ‘$<$’, etc.).
Table 2. Exploratory data analysis. Quantitative variables.
@@
revenue SD \$ 54,944.60 \$ 40,731.98 \$ 53,086.79
KS D-Test 0.09111Significant at the 0.01 level.
Mean 16.15 17.67 16.38
@@
dti_n SD 7.50 7.53 7.53
KS D-Test 0.09111Significant at the 0.01 level.
Mean \$ 13,799.25 \$ 15,111.44 \$ 13,999.66
@@
loan_amnt SD \$ 7,931.55 \$ 8,363.19 \$ 8,012.86
KS D-Test 0.08111Significant at the 0.01 level.
Mean 705.22 694.20 703.54
@@
fico_n SD 33.32 27.48 32.74
KS D-Test 0.15111Significant at the 0.01 level.
In Table 2 we show the quantitative variables and use the Kolmogorov-Smirnov test to compare the empirical cumulative probability distributions between the Default and Non-default loans. According to these results, defaulted loans are characterized by lower revenue, higher indebtedness (dti_n), higher requested amount (loan_amnt), and lower FICO (fico_n), being the differences significant at the 0.01 level.
Table 3. Exploratory data analysis. Categorical variables.
@@
purpose major purchase 3,062 2.57% 10.65% 568.47111Significant at the 0.01 level.
@@
Similarly, in Table 3 we show how the categories are distributed in each categorical variable and the default rate for each category. The addr_state is not included because it has 50 categories (one for each USA state). The table also shows whether there is a significant dependence between the target variable and the categorical variables at the 0.01 significance level. We can see that the dependence is significant for all the variables, including the addr_state not reported in the table (test value of 211.12).
In the home_ownership variable, the ‘OTHER’ category shows the highest risk (20.98%) but a small frequency (0.12%), while the ‘MORTGAGE’ category is the most frequent (51.05%) and the least risky (14.14%) one. In the emp_length variable, the category that denotes no information (‘NI’) has the highest risk (19.78%), but also the lowest frequency (4.13%). In general, the risk in employment length can be divided into two groups with similar default rates, over and below or equal to five years, and curiously the risk of the first group is slightly higher. The categories are not perfectly ordered. These facts support the inclusion of this variable as a categorical variable using one-hot encoding. Finally, the most frequent purpose is ‘debt consolidation’, constituting 57% of the loans, which has a default rate of 16.14%. Notably, the riskiest purpose is ‘small business’, with a 26.41% default rate. Conversely, ‘car’ loans demonstrate the lowest risk, with a mere 9.61% default rate. This striking divergence in default rates across diverse purposes underscores a significant variability in risk within the various loan purposes.
Table 4. Exploratory data analysis. Textual variable.
@@
Word count SD 46.62 48.78 46.96
KS D-Test 0.03111Significant at the 0.01 level.
Mean 66.70 66.31 66.64
@@
Readability SD 32.87 35.55 33.29
KS D-Test 0.02111Significant at the 0.01 level.
Mean 0.0964 0.0909 0.0956
@@
Polarity SD 0.1685 0.1699 0.1687
KS D-Test 0.04111Significant at the 0.01 level.
Mean 0.3193 0.3029 0.3168
@@
Subjectivity SD 0.2542 0.2595 0.2551
KS D-Test 0.04111Significant at the 0.01 level.
@@
5. Description of the methodology
This study seeks to assess the enhancement of default prediction by integrating textual descriptions of loans preprocessed through an LLM. Specifically, we employ transfer learning to enable the LLM to generate a score reflecting the probability of loan default, leveraging information derived from the loan’s textual description.
As our baseline, we employ a machine learning classifier that utilizes the complete set of variables available during the loan application process, encompassing both quantitative and categorical variables. Subsequently, we enhance the classifier by incorporating the LLM score—representing default risk inferred from processing the textual description with the LLM—as an additional input variable and proceed to evaluate the resulting differences. The data and code of our experiments are publicly available [^4] Link will be included in final version.
@@ 5. Description of the methodology
5.1. Tuning the classifier
The classification algorithm used was XGBoost (Chen and Guestrin, 2016), which is trained using a stratified k-fold cross-validation with k being equal to 5. A genetic algorithm (Holland, 1992) was employed to fine-tune the hyperparameters. The fitness value of each individual was the average balanced accuracy (BACC) in the 5 validation sets of the cross-validation. We use BACC as it accounts for the imbalanced nature of the dataset. In preliminary experiments, we also considered the area under the receiver operating characteristic (AUROC), which measures the model’s ability to discriminate between positive and negative examples regardless of the classification threshold chosen. However, we observed that the resulting XGBoost classifiers produced poor BACC values (similar to those from a naïve classifier that predicts the majority class) when using the standard 0.5 threshold to make the prediction. We also observed that XGBoost classifiers with extremely similar AUROC values produced very different results in terms of BACC. Thus, we decided to use the BACC measure as it resulted in classifiers with a more stable behavior.
@@ 5.1. Tuning the classifier
Table 5. Hyperparameters of XGBoost considered in the genetic optimization and their respective ranges.
@@
5.2. Generating a default score with BERT
@@
5.2.1. Transfer learning to produce the default score
The BERT model used [^5] Source: https://tfhub.dev/tensorflow/bert_en_uncased_L-12_H-768_A-12/4 has the following characteristics: L=12 hidden layers (i.e., Transformer encoder blocks), a hidden size of H=768, and A=12 attention heads (enabling its self-attention mechanism to operate concurrently in 12 distinct patterns). Consequently, the output generated by the BERT model are embeddings of size 768. For the transfer learning process, TensorFlow HUB was used due to its seamless integration with subsequent neural network layers.
To leverage the complex language representation of BERT, we decided to freeze the weights of all BERT hidden layers except the final one. Additionally, we augmented the BERT model by introducing supplementary layers post-embeddings to enrich its adaptability for our specific task and generate the default score. This architectural strategy aims to avoid catastrophic forgetting [^6] A common issue during retraining of pre-trained models that may lead to the deterioration of the original language representation., allowing the resulting model to specifically learn our task without compromising the previously acquired knowledge (Biesialska et al., 2020).
@@ 5.2.1. Transfer learning to produce the default score
5.2.2. Avoiding data leakage in cross-validation
@@
Figure 1. BERT training diagram.
5.2.3. Result of the transfer learning process
@@
Table 6. BERT optimal configuration and score by fold.
Neurons in
1st dense layer
Use 2nd dense layer
(128 neurons) [^1] Values ‘Yes’ or ‘No’ indicate whether the optimal configuration has or has not that layer.
Dropout Learning rate Loss  [^2] Weighted binary cross-entropy of the test set predictions.
Fold 0 512 Yes 20% 0.0001 0.1774
Fold 1 512 Yes 0% 0.00001 0.1793
Fold 2 128 Yes 20% 0.0001 0.1776
Fold 3 256 Yes 0% 0.001 0.1775
Fold 4 512 Yes 0% 0.001 0.1773
5.2.4. Inclusion of the BERT score in the classification model
@@
Figure 2. Diagram of the experiment architecture.
6. Analysis of the BERT score
Table 7. Loan descriptions with highest (top) and lowest (bottom) BERT score.
BERT score Real value
Description
@@ 6. Analysis of the BERT score
In this section, we analyze the BERT score generated. Table 7 shows the loan descriptions with the lowest and highest BERT scores—representing loans identified by BERT as having the highest and lowest default risks, respectively. At first sight, we can see that those with higher BERT scores correspond to poorly descriptive text containing typos (e.g., “pay of”, “everthing”, “alway”, “belive”) or not following grammar rules. Counter-intuitively, only one of these loans eventually defaulted. However, descriptions associated with lower BERT default predictions are more detailed, providing further insight into the purpose of the loan and emphasizing the borrower’s creditworthiness. Some even include numeric values that depict the financial condition of the borrower.
Figure 3. Default and non-default proportions by BERT score range.
In Figure 3 we analyze the proportion of default and non-default loans (Y axis) across the BERT score range (X axis) in bins of 0.01 parts per unit. In the bar, the blue part represents the percentage of Non-Default, while the orange part is the percentage of default observations [^10] The absence of a bar means that there are no loan descriptions with a BERT score within that range.. The figure shows that, in general, as the BERT score increases, the proportion of defaulted loans also increases. It is crucial to note that few observations fall outside the range of $[0.3,0.7]$, rendering the bars beyond this range less reliable. Nevertheless, as a general trend, a higher BERT score value tends to correspond to an elevated probability of default, which underscores the utility of the BERT score as a valuable risk indicator.
Table 8. Classification performance of BERT binarization at 0.5.
@@
These results show that the BERT model can be successfully fine-tuned to learn the task of predicting the final state of the loan using only the information provided by the borrower in the description field of the application form. In Section 7, we will further contextualize this result in light of the results provided by the XGBoost and different sets of variables.
6.1. Relationship between the BERT score and other variables
Table 9. Correlation coefficients of quantitative variables with the BERT score.
@@ 6.1. Relationship between the BERT score and other variables
Table 10. Kruskal-Wallis test of categorical variables with the BERT score.
@@
Table 11. Correlation coefficients of linguistic features with the BERT score.
@@
Finally, in Table 11, we investigate the potential relationship between the BERT score and various linguistic features automatically extracted from the text. We observe statistically significant correlations between the BERT score and all considered linguistic features. Notably, there is a relatively strong inverse correlation with both word count and subjectivity, implying that shorter and more objective texts may exhibit higher BERT scores. This observation, coupled with the correlations outlined in Table 9, suggests that BERT demonstrates a stronger association with linguistic aspects than with the numerical variables characterizing the loan application.
7. Classification results
7.1. Effect of including the BERT score
First, we evaluate the impact of incorporating the BERT score on classification performance. Table 12 compares the baseline experiment, featuring an optimized XGBoost with a genetic algorithm using quantitative and categorical variables, against the same approach augmented by the inclusion of the BERT score. A closer examination of the balanced metrics reveals a marginal enhancement in both BACC (0.6154 vs. 0.6187) and AUC (0.6575 vs. 0.6644), being the latter significant according to the DeLong test (Sun and Xu, 2014) at the 0.01 level. It is crucial to note that the Lending Club dataset exclusively consists of approved loans. This fact poses a substantial challenge to significantly enhance the outcomes in a loan granting model such as ours, since the loans included in the dataset were initially considered favorable by the platform.
Table 12. Classifier performance with and without the BERT score.

### v2

@@ 4 Description of the dataset
revenue Borrower’s self-declared annual income during registration.
dti_n Indebtedness ratio for obligations excluding mortgage. Monthly information.
loan_amnt Amount of credit requested by the borrower.
fico_n Credit bureau score. Defined between 300 and 850, reported by Fair Isaac Corporation as a summary risk measure based on historical credit information reported at the time of application.
@@
emp_length Employment length of the borrower categorized into 12 categories, including the no information category.
purpose Credit purpose category for the loan request.
home_ownership Homeownership status provided by the borrower.
addr_state Borrower’s residence state from the USA.
@@
desc Description of the credit request provided by the borrower.
In the dataset, the target variable suffers the usual class imbalance problem (only 15.27% of default), which will be considered in the design of the experiments. Table 1 shows the input variables of our granting model, which are explained below.
As for the quantitative variables, the Fair Isaac Corporation credit bureau (FICO) information in the original dataset is given by a minimum and maximum range of limits to which the borrower’s FICO belongs at loan origination. However, we average these two values to have a single indicator of the creditworthiness of potential borrowers resulting in our fico_n variable. For the case of the debt variable, dti_n is estimated from the original dataset variables as the ratio calculated from the total debts of the co-borrowers over the total debt obligation divided by the combined monthly income of the co-borrowers.
Regarding the categorical variables, we merged the categories ‘other’, ‘none’, and ‘any’ into a unified category labeled ‘other’ for the home_ownership variable. This decision was made due to a lack of clear differentiation among these options, coupled with their similar default percentages and their relatively low percentages of occurrences. The emp_length variable was treated as categorical rather than numerical since it includes categories for ‘no information’ and for ‘more than ten years’.
For the textual variable, we carried out an exhaustive work of text cleaning. First, we removed all those descriptions that contained the default description provided by Lending Club on its web form (“Tell your story. What is your loan for?”). Moreover, we removed the prefix “Borrower added on DD/MM/YYYY $>$ from the descriptions, as we did not want any temporal background on them. Finally, as these descriptions came from a web form, we replaced all HTML entities with their corresponding characters (e.g. ‘&amp;’ was substituted by ‘&’, ‘&lt;’ was substituted by ‘$<$’, etc.).
Table 2: Exploratory data analysis. Quantitative variables.
@@
SD \$ 54,944.60 \$ 40,731.98 \$ 53,086.79
KS D-Test 0.09*
dti_n Mean 16.15 17.67 16.38
@@
SD 7.50 7.53 7.53
KS D-Test 0.09*
loan_amnt Mean \$ 13,799.25 \$ 15,111.44 \$ 13,999.66
@@
SD \$ 7,931.55 \$ 8,363.19 \$ 8,012.86
KS D-Test 0.08*
fico_n Mean 705.22 694.20 703.54
@@
SD 33.32 27.48 32.74
KS D-Test 0.15*
* p-value less than 0.01.
Table 2 presents the quantitative variables and the results of the Kolmogorov-Smirnov test, which was used to compare the empirical cumulative distribution functions of Default and Non-default loans. According to these results, defaulted loans are characterized by lower revenue, higher debt-to-income ratio (dti_n), higher requested amount (loan_amnt), and lower FICO scores (fico_n), being the differences significant at the 0.01 level.
Table 3: Exploratory data analysis. Categorical variables.
@@
major purchase 3,062 2.57% 10.65%
@@
* p-value less than 0.01.
Similarly, Table 3 displays the distribution of categories within each categorical variable and the corresponding default rates. The addr_state variable is excluded due to its 50 categories, one for each U.S. state. The table also indicates whether there is a significant dependence between the target variable and the categorical variables at the 0.01 significance level. The results show significant dependence for all variables, including the addr_state variable not reported in the table (test value of 211.12).
In the home_ownership variable, the ‘OTHER’ category shows the highest risk (20.98%) but a small frequency (0.12%), while the ‘MORTGAGE’ category is the most frequent (51.05%) and the least risky (14.14%) one. In the emp_length variable, the category that denotes no information (‘NI’) has the highest risk (19.78%), but also the lowest frequency (4.13%). In general, employment length can be categorized into two groups with comparable default rates: those with employment lengths of five years or less and those with more than five years. Interestingly, the risk is slightly higher in the group with more than five years of employment. The categories are not perfectly ordered, which supports the use of one-hot encoding to treat this variable as categorical. Finally, the most frequent purpose is ‘debt consolidation’, constituting 57% of the loans, which has a default rate of 16.14%. Notably, the riskiest purpose is ‘small business’, with a 26.41% default rate. Conversely, ‘car’ loans demonstrate the lowest risk, with a mere 9.61% default rate. This striking divergence in default rates across diverse purposes underscores a significant variability in risk within the various loan purposes.
Table 4: Exploratory data analysis. Textual variable.
@@
SD 46.62 48.78 46.96
KS D-Test 0.03*
Readability Mean 66.70 66.31 66.64
@@
SD 32.87 35.55 33.29
KS D-Test 0.02*
Polarity Mean 0.0964 0.0909 0.0956
@@
SD 0.1685 0.1699 0.1687
KS D-Test 0.04*
Subjectivity Mean 0.3193 0.3029 0.3168
@@
SD 0.2542 0.2595 0.2551
KS D-Test 0.04*
* p-value less than 0.01.
@@
5 Description of the methodology
This study assesses the enhancement of default prediction by integrating textual descriptions of loans preprocessed through an LLM. Specifically, we employ transfer learning and fine-tuning to enable the LLM to generate a score reflecting the probability of loan default, leveraging information derived from the loan’s textual description.
As our baseline, we employ a machine learning classifier that utilizes the complete set of variables available during the loan application process, encompassing both quantitative and categorical variables. Subsequently, we enhance the classifier by incorporating the LLM score—representing default risk inferred from processing the textual description with the LLM—as an additional input variable and proceed to evaluate the resulting differences. The data and code of our experiments are publicly available [^4] Link of the code will be included in final version..
@@ 5 Description of the methodology
5.1 Tuning the classifier
The classification algorithm used was XGBoost [43], which is trained using a stratified k-fold cross-validation with k being equal to 5 and where the instances are shuffled to avoid potential ordering biases in the dataset. A genetic algorithm [44] was employed to fine-tune the hyperparameters. The fitness value of each individual was the average balanced accuracy (BACC) in the 5 validation sets of the cross-validation. We use BACC as it accounts for the imbalanced nature of the dataset. In preliminary experiments, we also considered the area under the receiver operating characteristic (AUROC), which measures the model’s ability to discriminate between positive and negative examples regardless of the classification threshold chosen. However, we observed that the resulting XGBoost classifiers produced poor BACC values (similar to those from a naïve classifier that predicts the majority class) when using the standard 0.5 threshold to make the prediction. We also observed that XGBoost classifiers with extremely similar AUROC values produced very different results in terms of BACC. Thus, we decided to use the BACC measure as it resulted in classifiers with a more stable behavior.
@@ 5.1 Tuning the classifier
Table 5: Hyperparameters of XGBoost considered in the genetic optimization and their respective ranges.
@@
5.2 Generating a default score with BERT
@@
5.2.1 Transfer learning to produce the default score
The BERT model we utilize [^5] Source: https://tfhub.dev/tensorflow/bert_en_uncased_L-12_H-768_A-12/4 is configured with L=12 hidden layers (i.e., Transformer encoder blocks), each with a size of H=768, and it employs A=12 attention heads. These attention heads enable the model’s self-attention mechanism to process inputs in 12 distinct patterns simultaneously. The output from BERT are embeddings of size 768. For incorporating this model, TensorFlow HUB was selected for its efficient integration with additional neural network layers.
As outlined in Section 3, during the transfer learning phase, we aim to exploit BERT’s advanced language understanding while minimizing the need to learn from scratch. To achieve this, we freeze the weights of all but the last hidden layer of BERT. This approach preserves the model’s pre-trained capabilities and prevents overfitting on our specific dataset, a phenomenon known as catastrophic forgetting [^6] A degradation issue during the retraining of pre-trained models, leading to loss of the original capabilities [45]..
Subsequently, in the fine-tuning stage, only the last BERT layer and the newly added layers are adjusted to better serve our specific task of generating a default score. These layers are optimized to enhance the model’s adaptability to our particular requirements, allowing slight parameter adjustments for improved task-specific performance, while maintaining the general language understanding gained from BERT’s initial pre-training. This strategy effectively balances specialized learning with the retention of valuable pre-trained knowledge.
@@ 5.2.1 Transfer learning to produce the default score
5.2.2 Avoiding data leakage in cross-validation
@@
Figure 1: BERT training diagram.
5.2.3 Result of the transfer learning process
@@
Table 6: BERT optimal configuration and score by fold.
Neurons in 1st dense layer Use 2nd dense layer (128 neurons)* Dropout Learning rate Loss** Fold 0 512 True 20% 0.0001 0.1774 Fold 1 512 True 0% 0.00001 0.1793 Fold 2 128 True 20% 0.0001 0.1776 Fold 3 256 True 0% 0.001 0.1775 Fold 4 512 True 0% 0.001 0.1773 * Values ‘True’ or ‘False’ indicate whether the optimal configuration has or has not that layer. ** Weighted binary cross-entropy of the test set predictions.
5.2.4 Inclusion of the BERT score in the classification model
@@
Figure 2: Diagram of the experiment architecture.
The fine-tuning of the BERT model was performed on a system equipped with an Intel Core i9-12900KS processor, an NVIDIA GeForce RTX 4090 graphics card (24GB), and 128 GB of RAM. In contrast, the inference process with the fine-tuned model, which involves generating the BERT score for a given loan description, can be conducted on a standard PC. This capability enhances the practicality of deploying our approach in real-world applications.
6 Analysis of the BERT score
6.1 Assessment of the BERT score as a credit risk score
Table 7: Loan descriptions with highest (top) and lowest (bottom) BERT score.
BERT score Real value Description
@@ 6.1 Assessment of the BERT score as a credit risk score
In this section, we evaluate the effectiveness of the BERT score in utilizing loan descriptions to predict default risk. Table 7 presents loan descriptions with the highest and lowest BERT scores, which indicate loans assessed by BERT as having the highest and lowest default risks, respectively. It appears that descriptions with higher BERT scores are often less informative, containing errors such as typos (e.g., “pay of”, “everthing”, “alway”, “belive”) or grammatical issues. Interestingly, only one of these loans ultimately defaulted. Conversely, loan descriptions with lower BERT scores tend to be more detailed and insightful, often including information about the loan’s purpose and the borrower’s creditworthiness, and sometimes even providing numeric data related to the borrower’s financial status.
Figure 3: Default and non-default proportions by BERT score range.
Figure 3 illustrates the distribution of default and non-default loans (Y axis) across the BERT score range (X axis) in bins of 0.01. In the bar chart, the blue segment represents the percentage of non-default loans, while the orange segment denotes the percentage of default loans [^10] Absence of a bar indicates that no loan descriptions fall within that BERT score range.. The figure reveals a general trend where higher BERT scores are associated with a higher proportion of defaulted loans. It is important to note that observations outside the BERT score range of $[0.3,0.7]$ are sparse, which makes the bars in these regions less reliable. Overall, the trend suggests that higher BERT scores are indicative of a greater likelihood of default, highlighting the BERT score’s usefulness as a risk assessment tool.
Table 8: Classification performance of BERT binarization at 0.5.
@@
These results demonstrate that the BERT model can be effectively fine-tuned to predict the final state of the loan using only the information provided by the borrower in the description field of the application form. In Section 7, we will contextualize these findings by comparing them with the results obtained from XGBoost and various sets of variables.
6.2 Assessment of the relationship between the BERT score and other variables
Table 9: Correlation coefficients of quantitative variables with the BERT score.
@@ 6.2 Assessment of the relationship between the BERT score and other variables
Table 10: Kruskal-Wallis test of categorical variables with the BERT score.
@@
Table 11: Correlation coefficients of linguistic features with the BERT score.
@@
Finally, Table 11 examines the potential relationship between the BERT score and various linguistic features automatically extracted from the text. We find statistically significant correlations between the BERT score and all linguistic features analyzed. Notably, there is a strong inverse correlation with both word count and subjectivity, indicating that shorter and more objective texts tend to have higher BERT scores. This finding, together with the correlations presented in Table 9, suggests that BERT is more closely related to linguistic features than to the numerical variables associated with loan applications.
7 Results of the LLM-enhanced granting model
7.1 Analysis of the classification performance
First, we evaluate the impact of incorporating the BERT score in the granting model. In our baseline experiment, we optimize XGBoost with a genetic algorithm using the quantitative and categorical variables typically used in granting models, while in the competing approach, we optimize it but include the BERT score as input variable. Table 12 shows the results of both approaches.
A closer examination of the balanced metrics reveals a marginal enhancement in both BACC (0.6154 vs. 0.6187) and AUC (0.6575 vs. 0.6644), the latter significant according to the DeLong test [46] at the 0.01 level. It is crucial to note that the Lending Club dataset exclusively consists of approved loans. This fact poses a substantial challenge to significantly enhance the outcomes in a loan granting model such as ours since the loans included in the dataset were initially considered favorable by the platform. Furthermore, in experiments not reported in the paper we used CatBoost [47] instead of XGBoost and obtained a similar BACC improvement.
Table 12: Classifier performance with and without the BERT score.